### v1

# MathMathgg

Abstract
We present an ab initio calculation of nuclear $\beta$-decay within the framework of nuclear lattice effective field theory (NLEFT), employing auxiliary-field quantum Monte Carlo methods to solve the nuclear many-body problem. Our approach combines next-to-next-to-leading order (NNLO) two- and three-body chiral interactions with one- and two-body axial current operators, all consistently derived in chiral effective field theory. Low-energy constants are determined exclusively from nucleon-nucleon scattering phase shifts and few-body observables for systems with $A\leq 3$. Using these interactions and transition operators, we perform two-channel Monte Carlo simulations to compute the $\beta$-decay matrix element for 6He, obtaining results in reasonable agreement with experimental measurements. To address the Monte Carlo sign problem, we implement a perturbative expansion around a leading-order Hamiltonian with approximate Wigner-SU(4) symmetry. This systematic approach provides a foundation for extending NLEFT simulations to precision studies of weak processes in medium-mass nuclei.
I Introduction
Nuclear $\beta$-decay is a fundamental process crucial to understanding the laws of nature. Historically, studies of $\beta$-decay have led to groundbreaking discoveries, including the existence of neutrinos Pauli (1978), the confirmation of parity violation Wu et al. (1957), and contributions to the unification of electroweak interactions Glashow (1961); Weinberg (1967); Salam (1968). Today, nuclear $\beta$-decay provides a sensitive probe for physics beyond the Standard Model. For a comprehensive review of ongoing and planned experimental programs, see Ref. González-Alonso et al. (2019) and references therein.
Besides its central role in particle physics, nuclear $\beta$-decay is also of great interest within the field of low-energy nuclear physics. Currently there remain several unresolved challenges in precisely describing the nuclear weak decay with the nucleonic degrees of freedom. For instance, traditional shell-model calculations tend to overestimate the reduced Gamow-Teller matrix element (RGTME), especially for medium-mass nuclei Brown and Wildenthal (1985); Martínez-Pinedo et al. (1996); Langanke et al. (1995). To reconcile this discrepancy, the axial charge $g_{A}$ in the axial current is often tuned to match experimental data Chou et al. (1993), leading to the well-known $g_{A}$ quenching problem. Moreover, the nuclear force binds nucleons into atomic nuclei, induces complicated many-body correlations, which requires tremendous amount of computing resources and advanced algorithms to solve the corresponding quantum many-body problems. Lastly, the external electroweak field interacts with the nucleus via not only one-body currents but also many-body currents, so a proper description of $\beta$-decays requires a consistent treatment of the electroweak currents together with the complicated nuclear chiral interactions.
Recent years have seen rapid development of $ab$ $initio$ nuclear many-body methods Carlson et al. (2015); Barrett et al. (2013); Hagen et al. (2014); Hergert et al. (2016); Lee (2009), which provide a consistent and systematically improvable framework for understanding the intricate interplays among nuclear forces, electroweak currents and quantum many-body correlations in $\beta$-decay processes. For nuclei of $s$ and $p$ shell, calculations based on the quantum Monte Carlo (QMC) method Pastore et al. (2018); King et al. (2020) and the no core shell model method Navrátil et al. (2007); Gysbers et al. (2019) have been performed systematically, with the latter applied to the study of the anomalous long lifetime of 14Maris et al. (2011). For heavier nuclei, the coupled-cluster method has been applied to investigating the impact of three-body forces and two-body currents on the Gamow-Teller transitions of 14C, 22O and 24Ekström et al. (2014), and it also provided the first $ab$ $inito$ calculation of 100Sn $\beta$-decay Gysbers et al. (2019). Besides, the valence-space in-medium similarity renormalization group method has made significant progress in studying $\beta$-decays of medium-mass nuclei Gysbers et al. (2019); Stroberg (2021); Yuan et al. (2025). These $ab$ $initio$ calculations have significantly improved our understanding of nuclear $\beta$-decays, yet also raised a lot of open questions. For instance, the QMC calculations struggle to reproduce the experimental $\beta$-decay rate for 8Li, 8B, and 8He King et al. (2020). Furthermore, the contribution of two-body currents is found to have opposite signs in Refs. King et al. (2020) and Gysbers et al. (2019). This theoretical inconsistency was further explored in Ref. Gnech et al. (2021), which also revealed that the one-pion-exchange (OPE) current is highly sensitive to the nuclear forces and regularization schemes.
As one of the advanced nuclear ab initio methods, the nuclear lattice effective field theory (NLEFT) Lee (2009); Lähde and Meißner (2019) implements the chiral interactions on a cubic spatial lattice and solves the corresponding discrete many-body Schrödinger equation using auxiliary-field Monte Carlo method. Recently the NLEFT has achieved remarkable progress in studying the properties of nuclear ground states and excited states Borasoy et al. (2007); Epelbaum et al. (2010a, b); Lähde et al. (2014); Lu et al. (2019, 2022); Elhatisari et al. (2024a); Epelbaum et al. (2014); Shen et al. (2023); Meißner et al. (2024); Shen et al. ; Zhang et al. ; Wang et al. , nuclear clustering Epelbaum et al. (2011, 2012, 2013); Elhatisari et al. (2016, 2017), scattering processes Bour et al. (2012); Elhatisari et al. (2015), thermal nuclear matter Lu et al. (2020); Ren et al. (2024); Ma et al. (2024a) , and hypernucleus Bour et al. (2015); Frame et al. (2020); Hildenbrand et al. (2024). In this area, recent years a significant amount of efforts was made to build high-precision lattice interactions and design algorithms that mitigate the Monte Carlo sign problem. For example, the wave function matching method Elhatisari et al. (2024a) and perturbative quantum Monte Carlo method Lu et al. (2022); Liu et al. were proposed to improve the accuracy of perturbative calculations around simple Hamiltonian without the sign problem. When combined with high-fidelity nuclear forces, these methods were able to reproduce the experimental binding energies and charge radii with systematically controllable statistical errors.
Attempts to study nuclear $\beta$-decay through NLEFT have just begun very recently. The lifetime of triton $\beta$-decay is calculated in Ref. Elhatisari et al. (2024b) using the Lanczos eigenvector method Lanczos (1950) to diagonalize the lattice Hamiltonian directly. However, for nuclei with mass number $A\geq 4$ the direct diagonalization is not possible due to the exponential increase of the computational resources. In this case we need the quantum Monte Carlo techniques which have polynomial scaling if the sign problem is well under control. Furthermore, the currents beyond leading order (LO) are omitted in Ref. Elhatisari et al. (2024b), leaving their effects undetected. In this regard, it is desirable to improve and apply the NLEFT Monte Carlo approach to investigating $\beta$-decays of light and medium-mass nuclei. The results would also provide an alternative benchmark for calculations with other ab initio methods.
In this work, we address this demand by presenting an in-depth study of nuclear $\beta$-decay within the NLEFT framework, focusing on the processes $\rightarrow{{}^{3}}$e and 6He $\rightarrow{{}^{6}}$i. The former serves as a benchmark for validating our methodology, while the latter illustrates its potential applicability to $\beta$-decays beyond few-body sector. We use next-to-next-to-leading order (NNLO) chiral interactions alongside nuclear axial currents up to next-to-next-to-next-to-leading order (N3LO), consistently derived from chiral EFT, to compute the RGTME. Since this is an exploratory study, our primary goal is to develop the necessary computational techniques and assess the reasonableness of our predictions. A more systematic study of $\beta$-decays in even heavier nuclei is left for the future.
This paper is organized as follows. The nuclear interaction and axial currents used in this work are described in Sec. II and III. Our calculation methods and techniques are briefly discussed in Sec. IV. Results for 3H and 6He $\beta$-decay are presented in Sec. V. More details are provided in the appendix.
II Interaction
II.1 Full Interaction
The full Hamiltonian used in this work is the NNLO chiral Hamiltonian discretized on a lattice with a lattice spacing $a=1.64$ fm,
$H=K+V_{\mathrm{OPE}}^{\Lambda_{\pi}}+V_{C_{\pi}}^{\Lambda_{\pi}}+V_{2\mathrm{N}}^{Q^{2}}+V_{\mathrm{IB}}+V^{Q^{3}}_{3\mathrm{N}}.$ (1)
Here, $K$ is kinetic term, and the nucleon mass is taken as $M_{N}=938.92$ MeV. The OPE potential is locally regularized in momentum space, following the prescription of Ref. Reinert et al. (2018),
$V_{\mathrm{OPE}}^{\Lambda_{\pi}}=-\frac{\tilde{g}_{A}^{2}}{4f_{\pi}^{2}}(\bm{\tau}_{1}\cdot\bm{\tau}_{2})\frac{\bm{(\sigma}_{1}\cdot\bm{q})(\bm{\sigma}_{2}\cdot\bm{q})}{\bm{q}^{2}+M_{\pi}^{2}}e^{-\frac{\bm{q}^{2}+M_{\pi}^{2}}{\Lambda_{\pi}^{2}}},$ (2)
$V_{C_{\pi}}^{\Lambda_{\pi}}=-\frac{\tilde{g}_{A}^{2}C_{\pi}}{4f_{\pi}^{2}}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})(\bm{\tau}_{1}\cdot\bm{\tau}_{2}),$ (3)
where
$C_{\pi}=\frac{\Lambda_{\pi}(\Lambda_{\pi}^{2}-2M_{\pi}^{2})+2\sqrt{\pi}M_{\pi}^{3}\exp\left(\frac{M_{\pi}^{2}}{\Lambda_{\pi}^{2}}\right)\mathrm{erfc}\left(\frac{M_{\pi}}{\Lambda_{\pi}}\right)}{3\Lambda_{\pi}^{3}}.$ (4)
The numerical values of the parameters used are $\tilde{g}_{A}=1.287$ for the nucleon axial charge, $f_{\pi}=92.2$ MeV for the pion decay constant, $M_{\pi}=134.98$ MeV for the pion mass, and $\Lambda_{\pi}=300$ MeV for the local regulator. The axial charge $\tilde{g}_{A}$ incorporates the Goldberger-Treiman discrepancy Arndt et al. (1994) and differs slightly from value $g_{A}=1.2723(23)$ Tanabashi et al. (2018) in the axial current. To distinguish between them, we use a tilde notation. The term in Eq. (3) serves as a counterterm to remove the short-distance singularity in the OPE potential. It is worth noting that the characteristic momentum scale in this work is around 100 MeV, considerably lower than twice the pion mass. As a result, the two-pion exchange (TPE) potential, which first appears at NLO, cannot be distinguished from contact terms Epelbaum et al. (2009).
The two-nucleon (2N) short-range potential, $V^{Q^{2}}_{2\mathrm{N}}$, includes two contact terms at leading order (LO) and seven additional contact terms at next-to-leading order (NLO),
$V_{2\mathrm{N}}^{Q^{2}}$ $=$ $B_{1}+B_{2}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})+C_{1}q^{2}+C_{2}q^{2}(\bm{\tau}_{1}\cdot\bm{\tau}_{2})$ (5)
$+$ $C_{3}q^{2}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})+C_{4}q^{2}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})(\bm{\tau}_{1}\cdot\bm{\tau}_{2})$
$+$ $C_{5}(\bm{\sigma}_{1}\cdot\bm{q})(\bm{\sigma}_{2}\cdot\bm{q})+C_{6}(\bm{\sigma}_{1}\cdot\bm{q})(\bm{\sigma}_{2}\cdot\bm{q})(\bm{\tau}_{1}\cdot\bm{\tau}_{2})$
$+$ $\frac{i}{2}C_{7}(\bm{q}\times\bm{k})\cdot(\bm{\sigma}_{1}+\bm{\sigma}_{2}),$
where $B_{i}$ and $C_{i}$ are low-energy constants (LECs) fitted to the Nijmegen neutron-proton ($n$-$p$) scattering phase shifts Stoks et al. (1993). All contact terms are regularized using a nonlocal Gaussian regulator in momentum space, following Ref. Lu and Deng . Specifically, for the matrix element $V(\bm{p}_{1},\bm{p}_{1}^{\prime},\bm{p}_{2},\bm{p}_{2}^{\prime})$, where $\bm{p}_{i}$ and $\bm{p}^{\prime}_{i}$ are the incoming and outgoing momenta of the $i$-th nucleon, we apply a non-local momentum regulator
$f_{2\mathrm{N}}=\prod_{i=1}^{2}\exp\left(-\frac{p_{i}^{6}+p_{i}^{\prime 6}}{2\Lambda^{6}}\right).$ (6)
In the center-of-mass frame, Eq. (6) reduces to the nonlocal regulator commonly used in Ref.  Entem and Machleidt (2003).
The isospin-breaking term, $V_{\mathrm{IB}}$, first appears at NLO. Following the convention of Ref. Epelbaum et al. (2010a), we separate it into three components: the long-range Coulomb term,
$V_{\mathrm{IB}}^{\mathrm{Cou}}=\frac{\alpha_{\mathrm{EM}}}{\mathrm{max}(|\bm{r}_{1}-\bm{r}_{2}|,0.5)}\left(\frac{1+\tau_{1,z}}{2}\right)\left(\frac{1+\tau_{2,z}}{2}\right),$ (7)
the short-range proton-proton ($p$-$p$) contact term,
$V_{pp}=C_{pp}\delta(\bm{r}_{1}-\bm{r}_{2})\left(\frac{1+\tau_{1,z}}{2}\right)\left(\frac{1+\tau_{2,z}}{2}\right),$ (8)
and the short-range neutron-neutron ($n$-$n$) contact term,
$V_{nn}=C_{nn}\delta(\bm{r}_{1}-\bm{r}_{2})\left(\frac{1-\tau_{1,z}}{2}\right)\left(\frac{1-\tau_{2,z}}{2}\right).$ (9)
These isospin-breaking terms are essential for reproducing the 3H-3He mass splitting Epelbaum et al. (2010b). The 2N LECs $C_{pp}$ and $C_{nn}$ in Eqs. (8)-(9) are fitted to experimental $p$-$p$ phase shifts, as well as $n$-$n$ scattering lengths and effective ranges Gonzalez Trotter et al. (2006); Miller et al. (1990). Details on determining the LECs are provided in Appendix A.
Finally, for the three-nucleon (3N) potential $V_{3\mathrm{N}}^{Q^{3}}$, which first appears at NNLO, we include the OPE term
$V_{3\mathrm{N}}^{\mathrm{OPE}}$ $=$ $-\frac{g_{A}}{8f^{2}_{\pi}}\frac{c_{D}}{f_{\pi}^{2}\Lambda_{\chi}}\sum_{C(i,j,k)}\frac{(\bm{\sigma}_{i}\cdot\bm{q}_{i})(\bm{\sigma}_{j}\cdot\bm{q}_{i})}{\bm{q}_{i}^{2}+M_{\pi}^{2}}(\bm{\tau}_{i}\cdot\bm{\tau}_{j}),$
and the 3N contact term
$V_{3\mathrm{N}}^{\mathrm{CT}}=\frac{c_{E}}{2f_{\pi}^{4}\Lambda_{\chi}}\sum_{C(i,j,k)}\bm{\tau}_{i}\cdot\bm{\tau}_{j}.$ (11)
Here, $\Lambda_{\chi}=700$ MeV is the chiral symmetry breaking scale. The 3N regulator
$f_{3\mathrm{N}}=\prod_{i=1}^{3}\exp\left(-\frac{p_{i}^{6}+p_{i}^{\prime 6}}{2\Lambda^{6}}\right)$ (12)
is employed to regularize $V_{3\mathrm{N}}^{\mathrm{CT}}$. We have also tested the 3N TPE potential but found its contribution negligible, so we neglect it in this work. The LECs $c_{D}$ and $c_{E}$ are determined by fitting to the experimental 3H binding energy and the RGTME of 3H $\beta$-decay process simultaneously. Details are given in Sec. V.
II.2 Non-perturbative Interaction
The full Hamiltonian in Eq. (1) includes complex spin-isospin-dependent terms, which pose significant challenges for lattice calculations. One major issue is that these terms can induce severe sign problems in fermionic systems, such as nucleons, in Monte Carlo simulations Lähde and Meißner (2019). To mitigate this, one can construct a simplified Hamiltonian $H_{0}$ which captures the essential physics but with much milder sign problems, and solve the full Hamiltonian $H$ perturbatively Lu et al. (2022). Details of the perturbative method will be provided in Sec. IV.1. In this work, we consider the following non-perturbative Hamiltonian:
$H_{0}$ $=$ $K+\frac{C_{2}}{2!}\sum_{\bm{n}}:\tilde{\rho}^{2}(\bm{n}):+\frac{C_{3}}{3!}\sum_{\bm{n}}:\tilde{\rho}^{3}(\bm{n}):$ (13)
$+$ $\frac{C_{I}}{2!}\sum_{I,\bm{n}}:\tilde{\rho}^{2}_{I}(\bm{n}):+V_{\mathrm{OPE}}^{\Lambda_{\pi}^{\prime}}.$
Here, :: represents normal ordering and $\tilde{\rho}$ is the SU(4)-symmetric density operator smeared both locally and nonlocally Elhatisari et al. (2017):
$\tilde{\rho}(\bm{n})$ $=$ $\sum_{s,i}\tilde{a}^{\dagger}_{s,i}(\bm{n})\tilde{a}_{s,i}(\bm{n})$ (14)
$+$ $s_{\mathrm{L}}\sum_{|\bm{n}-\bm{n}^{\prime}|=1}\sum_{s,i}\tilde{a}^{\dagger}_{s,i}(\bm{n}^{\prime})\tilde{a}_{s,i}(\bm{n}^{\prime}),$
where $s$ and $i$ are the spin and isospin index, respectively. The parameter $s_{\mathrm{L}}$ controls the range of the local part of the interaction. The smeared creation operator is defined as
$\tilde{a}^{\dagger}_{s,i}(\bm{n})=a^{\dagger}_{s,i}(\bm{n})+s_{\mathrm{NL}}\sum_{|\bm{n}-\bm{n}^{\prime}|=1}\tilde{a}^{\dagger}_{s,i}(\bm{n}^{\prime}),$ (15)
with $s_{\mathrm{NL}}$ controlling the range of the nonlocal part of the interaction. The definition of the isospin-dependent density operator $\tilde{\rho}_{I}$ is similar to $\tilde{\rho}$:
$\tilde{\rho}_{I}(\bm{n})$ $=$ $\sum_{s}\sum_{i,i^{\prime}}\tilde{a}^{\dagger}_{s,i}(\bm{n})[\tau_{I}]_{i,i^{\prime}}\tilde{a}_{s,i^{\prime}}(\bm{n})$
$+$ $s_{\mathrm{L}}\sum_{|\bm{n}-\bm{n}^{\prime}|=1}\sum_{s}\sum_{i,i^{\prime}}\tilde{a}^{\dagger}_{s^{\prime},i}(\bm{n}^{\prime})[\tau_{I}]_{i,i^{\prime}}\tilde{a}_{s^{\prime},i^{\prime}}(\bm{n}^{\prime}),$
with $\tau_{I}$ the $I$-th Pauli matrix .
The Hamiltonian in Eq. (13) consists of two parts, written separately for clarity. The first line contains terms that respect Wigner’s SU(4) symmetry Wigner (1937), an approximate symmetry of low-energy nuclear interactions. The inclusion of this symmetry helps suppress sign problems in lattice Monte Carlo calculations Chen et al. (2004). Recently, the SU(4)-symmetric Hamiltonian has been applied to the study of the properties of light and medium-mass nuclei as well as the equation of state of symmetric nuclear matter and neutron matter Lu et al. (2019), the intrinsic clustering geometry of 12C’s low-lying states Shen et al. (2023), the thermodynamics of nuclear systems Lu et al. (2020), the clustering effects in hot, dilute nuclear matter Ren et al. (2024), and the alpha-particle monopole transition form factor Meißner et al. (2024). The second line of Eq. (13) introduces an isospin-dependent contact term and a OPE term, both of which explicitly break the SU(4) symmetry. These additional terms are introduced to improve the overall description of $p$-shell nuclei Wiringa and Pieper (2002) , such as 6He and 6Li studied in this work. To suppress the sign problem caused by the tensor force, we adopt a lower cutoff of $\Lambda_{\pi}^{\prime}=180$ MeV for the OPE term in Eq. (13), compared to the higher cutoff of $\Lambda_{\pi}=300$ MeV used in the full Hamiltonian in Eq. (1).
To determine the unknown parameters in Eq. (13), we perform a combined fit to $n$-$p$ scattering phase shifts in two $S$-wave channels, along with the ground-state binding energies of 4He and 6He. The inclusion of the latter two observables is motivated by the fact that 6He is a weakly bound halo nucleus Tanihata (1996); Jensen et al. (2004); Zhukov et al. (1993), with its binding energy very close to that of 4He:
$E(^{4}\mathrm{He})=-28.30\mathrm{MeV},\ E(^{6}\mathrm{He})=-29.27\mathrm{MeV}.$ (17)
Therefore, we carefully tune $H_{0}$ to ensure that 6He is bound in our calculation. The fitted parameter values are presented in Table 1.
$s_{\mathrm{L}}$ $s_{\mathrm{NL}}$ $C_{2}(\mathrm{MeV}^{-2})$ $C_{I}(\mathrm{MeV}^{-2})$ $C_{3}(\mathrm{MeV}^{-5})$
0.109 0.122 -6.41$\times$10-6 9.20$\times$10-7 -9.65$\times$10-13
Table 1: Values for all the parameters in Eq. (13).
III Axial Current
The transition operator for the RGTME is related to the nuclear axial current. We employ one- and two-body $\Delta$-less chiral EFT axial currents derived from the unitary transformation method of the Bonn group Krebs et al. (2017) and the time-ordered perturbation method of the JLab-Pisa group Baroni et al. (2016a) . These two approaches yield consistent tree-level results up to N3LO ($Q^{0}$) but exhibit some differences at N4LO ($Q$) due to loop contributions Baroni et al. (2016b); Krebs et al. (2017, 2020). Moreover, Ref. Krebs (2020) recently highlighted that the absence of consistent regularization for axial-vector current operators leads to a violation of chiral symmetry in the chiral limit at N4LO. For these reasons, we restrict our analysis to tree-level axial currents up to N3LO and neglect higher-order contributions.
Figure 1: Diagrams illustrating the contributions to the axial current up to N3LO used in this work. Nucleons, pions and the external field are denoted by the solid , dashed, and wavy lines respectively. Panel (a) denotes the LO current while Panel (b) is the relativistic correction at NNLO. Panel (c) and Panel (d) are the OPE current and the contact current at N3LO, respectively.
Since the momentum carried by the current is very small in $\beta$-decay processes, we work in the limit of vanishing current momenta to simplify calculations. The explicit momentum-space (MS) expressions of the axial currents under this limit are provided in Ref. Baroni et al. (2016b), and a diagrammatic representation of the currents is shown in Fig. 1. For two-body currents at N3LO, we regularize them using the regulator $f_{2\mathrm{N}}$ introduced in Eq. (5). The N3LO OPE current involves the two $N$-$\pi$ coupling constants, $c_{3}$ and $c_{4}$, for which we adopt the values $c_{3}=-5.61$ GeV-1 and $c_{4}=4.26$ GeV-1 from Ref. Hoferichter et al. (2015).
Here, we highlight a subtle difference between our calculation on a discretized cubic lattice and previous studies in the continuum Park et al. (2003); Baroni et al. (2016b, 2018); Gnech et al. (2021); King et al. (2020). In the latter case, the MS expression of the current is first transformed into configuration space (CS) analytically, i.e., $\bm{J}_{5}^{\mathrm{MS}}\rightarrow\bm{J}_{5}^{\mathrm{CS}}$. In this step, some parts of the OPE current, $\bm{J}_{5}^{\mathrm{MS}}(\mathrm{OPE})$, are absorbed into the definition of the contact current, $\bm{J}_{5}^{\mathrm{CS}}(\mathrm{CT})$, using Fermi-Dirac statistics Park et al. (2003). As a result, if the MS expression of the contact term is
$\bm{J}^{\mathrm{MS}}_{5,a}(\mathrm{CT})=z_{0}(\bm{\tau}_{1}\times\bm{\tau}_{2})_{a}\bm{\sigma}_{1}\times\bm{\sigma}_{2},$ (18)
then its CS expression takes the form
$\bm{J}^{\mathrm{CS}}_{5,a}(\mathrm{CT})=z^{\prime}_{0}(\bm{\tau}_{1}\times\bm{\tau}_{2})_{a}\bm{\sigma}_{1}\times\bm{\sigma}_{2}.$ (19)
The two coefficients $z_{0}$ and $z_{0}^{\prime}$ are related to each other through
$z_{0}^{\prime}\propto z_{0}+\frac{1}{6}c_{3}+\frac{1}{3}c_{4}+\frac{1}{12M_{N}},$ (20)
with the last three terms $c_{3}/6$, $c_{4}/3$, and $1/(12M_{N})$ originally belonging to $\bm{J}_{5}^{\mathrm{MS}}(\mathrm{OPE})$. However, in our lattice calculation, since both the momenta and the coordinates are discretized, the transformation from MS to CS can be efficiently handled numerically using fast Fourier transform, and these is no need to perform any analytical calculation. Therefore, we do not split the OPE current into two parts, and the MS and CS currents remain in one-to-one correspondence, eliminating any ambiguity. Finally, we note that the coefficient $z_{0}$ in Eq. (18) is related to the three-body LEC $c_{D}$ via Baroni et al. (2018)
$z_{0}=-\frac{c_{D}}{8\Lambda_{\chi}f_{\pi}^{2}},$ (21)
and we take $\Lambda_{\chi}=700$ MeV for the consistency with Eq (11).
IV Computational Methods
In this section, we introduce two methods used in this work, the Monte Carlo method and the sparse-matrix diagonalization method, to calculate the RGTME
$\mathrm{GT}=\frac{\sqrt{2J_{f}+1}}{g_{A}}\frac{\langle J_{f}M_{f}|\mathcal{J}|J_{i}M_{i}\rangle}{\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}.$ (22)
Here, $|J_{i}M_{i}\rangle$ and $|J_{f}M_{f}\rangle$ represent the eigenstates of the full Hamiltonian $H$ for the initial and final nucleus, respectively, with $J$ the nuclear spin and $M$ the magnetic quantum number. The weak transition operator $\mathcal{J}$ is defined as
$\mathcal{J}=\frac{1}{\sqrt{2}}(J_{5,+}^{x}+iJ_{5,+}^{y}),$ (23)
where $J_{5,+}^{x/y}$ is the $x/y$-component of the charge-raising axial current. Since $\mathcal{J}$ is a rank-one spherical tensor carrying the magnetic quantum number $m_{\mathcal{J}}=1$, the Clebsch-Gordan coefficient $\langle J_{i}M_{i},11|J_{f}M_{f}\rangle$ appears in the denominator for the extraction of the reduced matrix element. For the processes studied in this work, we have $J_{i}=J_{f}=\frac{1}{2}$ for $\rightarrow^{3}$e $\beta$-decay and $J_{i}=0,J_{f}=1$ for $\rightarrow^{6}$i $\beta$-decay, and in our calculation, we set $M_{i}=-1/2$, $M_{f}=1/2$ for the former transition and $M_{i}=0$, $M_{f}=1$ for the latter.
For convenience in later analysis, we decompose $\mathcal{J}$ into different orders according to Fig. 1:
$\mathcal{J}=\mathcal{J}_{\mathrm{LO}}+\mathcal{J}_{\mathrm{N}^{2}\mathrm{LO}}+\mathcal{J}_{\mathrm{N}^{3}\mathrm{LO}}.$ (24)
Accordingly, the RGTME can be expanded order by order as
$\mathrm{GT}=\mathrm{GT}_{\mathrm{LO}}+\mathrm{GT}_{\mathrm{N}^{2}\mathrm{LO}}+\mathrm{GT}_{\mathrm{N}^{3}\mathrm{LO}}.$ (25)
IV.1 Monte Carlo Method
The Monte Carlo method is a crucial tool for simulating nuclear systems in NLEFT. In this section, we discuss its application to our work.
We begin by introducing the full transfer matrix
$M=:e^{-a_{t}H}:$ (26)
for the full Hamiltonian $H$ defined in Eq. (1), where $a_{t}$ is the temporal lattice spacing. To facilitate perturbative calculations that will be introduced below, we also define a non-perturbative transfer matrix
$M_{0}=:e^{-a_{t}H_{0}}:,$ (27)
corresponding to the non-perturbative Hamiltonian $H_{0}$ defined in Eq. (13).
Next, we prepare the trial states $|\Phi_{i/f}\rangle$, which share the same quantum numbers as $|J_{i/f}M_{i/f}\rangle$. These states are evolved along Euclidean time using the transfer matrix $M$,
$|\Phi^{t/2}_{i/f}\rangle\equiv M^{N_{t}/2}|\Phi_{i/f}\rangle,$ (28)
where $N_{t}$ is the number of evolution steps and $t=N_{t}a_{t}$ represents the evolution time. Since both processes studied here, i.e., ${}^{3}\mathrm{H}\rightarrow^{3}\mathrm{He}$ and ${}^{6}\mathrm{He}\rightarrow$ ${}^{6}\mathrm{Li}$, only involve ground-state nuclei, $t$ is taken as large as possible to suppress excited-state contamination. To accelerate the convergence, the trial states $|\Psi_{i/f}\rangle$ are constructed based on the the nuclear shell model Mayer (1948); Heyde (1994).
The RGTME, given in Eq. (22), can be extracted from the ratio
$\mathrm{GT}(t)=\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi^{t/2}_{f}|\mathcal{J}|\Phi^{t/2}_{i}\rangle}{\sqrt{\langle\Phi^{t/2}_{i}|\Phi^{t/2}_{i}\rangle\langle\Phi^{t/2}_{f}|\Phi^{t/2}_{f}\rangle}}$ (29)
at large $t$. To evaluate it, auxiliary field transformations are applied to both the numerator and denominator, reducing the problem to a multi-dimensional integral that can be handled using Monte Carlo methods. A detailed discussion of this process is provided in Appendix C. However, as is discussed in Section II.2, a direct auxiliary field transformation on the full transfer matrix $M$ leads to severe sign problems, making the results highly noisy.
To address this issue, we employ perturbation theory Lu et al. (2022). Specifically, we decompose the transfer matrix into
$M=M_{0}+\Delta M,$ (30)
treating $\Delta M$ as a perturbation. This allows us to expand Eq. (29) as a perturbative series:
$\mathrm{GT}=\sum_{n=0}^{\infty}\mathrm{GT}^{(n)}_{\mathrm{LO}}+\sum_{n=0}^{\infty}\mathrm{GT}^{(n)}_{\mathrm{N}^{2}\mathrm{LO}}+\sum_{n=0}^{\infty}\mathrm{GT}^{(n)}_{\mathrm{N}^{3}\mathrm{LO}},$ (31)
with the variable $t$ suppressed for brevity. Here, $n$ denotes the order of expansion in $\Delta M$. The contributions from different orders of the current are written as LO, N2LO and N3LO, respectively. For the LO current, we retain the first two terms in the perturbative series and neglect higher-order corrections, namely,
$\mathrm{GT}_{\mathrm{LO}}\approx\mathrm{GT}^{(0)}_{\mathrm{LO}}+\mathrm{GT}^{(1)}_{\mathrm{LO}}.$ (32)
For currents beyond LO, we keep only the first term,
$\mathrm{GT}_{\mathrm{N}^{2}\mathrm{LO}/\mathrm{N}^{3}\mathrm{LO}}\approx\mathrm{GT}^{(0)}_{\mathrm{N}^{2}\mathrm{LO}/\mathrm{N}^{3}\mathrm{LO}}.$ (33)
Thus, Eq. (31) is approximated as
$\mathrm{GT}$$\approx$ $\sum_{n=0}^{1}\mathrm{GT}^{(n)}_{\mathrm{LO}}+\mathrm{GT}^{(0)}_{\mathrm{N}^{2}\mathrm{LO}}+\mathrm{GT}^{(0)}_{\mathrm{N}^{3}\mathrm{LO}}.$ (34)
The explicit expressions for the terms on the right-hand side of Eq. (34) are provided in Appendix B. They are computed using Monte Carlo methods. Since the system now evolves under $H_{0}$ that has an approximate Wigner-SU(4) symmetry instead of the full Hamiltonian $H$, the sign problem is significantly alleviated.
When designing the algorithm, we note that any $\beta$-decay process involves two different nuclei, making this essentially a two-channel problem. To control statistical errors, an appropriate probability distribution function is required to update the auxiliary fields during the Markov process. We leave relevant details in Appendix C.
IV.2 Lanczos Eigenvector Method
For the $\rightarrow{{}^{3}}$e process, the dimension of the Hilbert space is not too large and a rigorous calculation is possible. Therefore, in addition to the Monte Carlo approach, we employ the Lanczos eigenvector technique as an alternative method Epelbaum et al. (2010b). Specifically, we directly diagonalize the full Hamiltonian $H$ to obtain the ground-state wave functions of 3H and 3He. These wave functions are then used as input to compute the RGTME following Eq. (22). We make the calculation on a box with its length $L\approx 20$ fm, which is sufficiently large to neglect finite-volume effects Elhatisari et al. (2024b).
The main advantage of the Lanczos method is its accuracy. Unlike Monte Carlo simulations, it is free from statistical errors and systematic errors from the perturbative method. Additionally, due to its precision, this method enables a precise determination of the three-body LECs $c_{D}$ and $c_{E}$ from 3H $\beta$-decay. Specifically, we fit them to the experimental binding energy and half-life of 3H, closely following the procedure in Refs. Gazit et al. (2009); Baroni et al. (2016b, 2018). The details are presented in the next section.
V Results
V.1 3H $\beta$-decay
The details of 3H $\beta$-decay are discussed in this subsection. An essential step here is to convert the half life, $t_{1/2}$, into the RGTME. We make use of the formula Simpson (1987); Hardy et al. (1990); Schiavilla et al. (1998)
$(1+\delta_{R})t_{1/2}f_{V}=\frac{2\pi^{3}\mathrm{ln}2/[m_{e}^{5}V_{ud}^{2}G_{F}^{2}(1+\Delta_{R}^{V})]}{\mathrm{F}^{2}+(f_{A}/f_{V})g^{2}_{A}\mathrm{GT}^{2}},$ (35)
which expresses $t_{1/2}$ in terms of the Fermi matrix element
$\mathrm{F}=\sum_{n=1}^{3}\langle^{3}\mathrm{He}||\tau_{n,+}||^{3}\mathrm{H}\rangle$ (36)
and the RGTME. In Eq. (35), $m_{e}=0.511$ MeV is the electron mass, $V_{ud}=0.9737(3)$ is the up-down quark-mixing element of the Cabibbo-Kobayashi-Maskawa matrix Hardy and Towner (2020a), $G_{F}=1.1664\times 10^{-5}$ GeV-1 is the Fermi constant Tishchenko et al. (2013), $f_{V}=2.8355\times 10^{-6}$ and $f_{A}=2.8505\times 10^{-6}$ are Fermi functions Simpson (1987), $\delta_{R}=1.9\%$ is the outer radiative correction of 3H $\beta$-decay Raman et al. (1978), and $\Delta_{R}^{V}=0.02467(22)$ is the inner radiative correction Seng et al. (2018). Our lattice result of the Fermi matrix element is
$\mathrm{F}=0.9996,$ (37)
combined with the measured comparative half-life, $(1+\delta_{R})t_{1/2}f_{V}=1134.6(31)$Simpson (1987), determines the value of the RGTME:
$\mathrm{GT}_{\mathrm{exp}}=1.6473(41).$ (38)
The subscript “$\mathrm{exp}$” means that the result is derived from experimental data, although the lattice prediction of the Fermi matrix element is used as an input. We incorporate the errors in $g_{A},V_{ud}$ and $\Delta_{R}^{V}$, therefore the uncertainty above is larger than the one in Ref. Elhatisari et al. (2024b).
Now we show the determination of $c_{D}$ and $c_{E}$ using the Lanczos eigenvector method. In Fig. 2, $c_{E}$ is plotted against $c_{D}$ by fitting to the experimental binding energy $E(^{3}$)=8.482)=-8.482) = - 8.482 MeV. Similar to previous studies, we observe a nearly linear relation between $c_{D}$ and $c_{E}$ Gazit et al. (2009); Baroni et al. (2016b, 2018). Then we calculate the lattice RGTME, $\mathrm{GT}_{\mathrm{lat}}$, and plot its ratio with respect to $\mathrm{GT}_{\mathrm{exp}}$ in Fig. 3, as a function of $c_{D}$. All the points in Fig. 3 share the same values of $(c_{D},c_{E})$ with the ones in Fig. 2. Again, GTth/GTexp is a nearly linear function of $c_{D}$, consistent with Ref. Baroni et al. (2018). The values of $c_{D}$ and $c_{E}$ are determined as
$c_{D}=1.01(16),\quad c_{E}=0.81(1).$ (39)
Figure 2: The $c_{E}$-$c_{D}$ trajectories obtained by fitting the experimental 3H binding energy 8.482 MeV.
Figure 3: The ratio GTlat/GTexp as a function of $c_{D}$. The gray band shows the experimental uncertainty of GTexp.
LO NNLO N3LO(OPE) N3LO(CT)
LE 1.6747 $-0.0083$ $-0.0480$ 0.0289
MC 1.6652(121) $-0.0071(14)$0.0509(37) 0.0273(22)
Table 2: Contributions to RGTME from currents at different orders for $c_{D}=$1.01, $c_{E}$=0.81. The upper and lower lines are the results of the Lanczos eigenvector method(LE) and the Monte Carlo (MC) method, respectively. The value in the bracket is the statistical error.
Figure 4: RGTME of 6He $\beta$-decay versus the Euclidean evolution time $t$. In the upper panel, the black points represent the contribution of the LO axial current, the red points denote the full RGTME $\mathrm{GT}$ and the red band is the fitting result. The experimental value of $\mathrm{GT}$ is plotted as the dashed purple line. The lower panel shows the net contributions from currents beyond LO.
LO NNLO N3LO(OPE) N3LO(CT) Total-LO Total Exp Knecht et al. (2012)
RGTME 2.289(19) -0.024(2) -0.104(1) 0.032(1) -0.096(2) 2.192(20) 2.161(4)
Table 3: Contributions to RGTME of 6He $\beta$-decay from currents at different orders. Columns labeled with LO, NNLO, N3LO(OPE) and N3LO(CT) correspond to the diagrams illustrated in panel (a),(b),(c) and (d) of Fig. 1, respectively. The full result is given in the column labeled “Total”, while results including only corrections beyond LO are listed under “Total-LO”. The value in the bracket is the statistical uncertainty. The experimental value in the last column is cited from Ref. Knecht et al. (2012).
To evaluate the contributions from currents at different orders , we calculate the corresponding RGTMEs and present the results in Table 2. It is evident that the LO current provides the dominant contribution, while corrections from higher-order currents remain below 5%. To validate the two computational approaches described in Sec. IV, we also performed Monte Carlo calculations for 3H $\beta$-decay. The results of both approaches, shown in the second and third rows of Table 2, exhibit good agreement.
[…]
V.2 6He $\beta$-decay
We first present the energies and charge radii of 6He and 6Li in Table. 4. The energy is computed up to second-order perturbation theory Lu et al. (2022), leading to relatively large uncertainties. The charge radius, on the other hand, is calculated up to the first order. Overall, our results show satisfactory agreement with experimental data.
Nuclei($J^{P}$) 6He(0+) 6Li(1+)
$E$/MeV $R_{\mathrm{ch}}$/fm $E$/MeV $R_{\mathrm{ch}}$/fm
Lat -28.80(1.21) 1.988(32) -31.88(93) 2.534(18)
Exp -29.27 2.068(11) -31.99 2.589(39)
Table 4: Ground state energies and the charge radii of 6He and 6Li calculated on the lattice compared to experiment. The third line is the lattice prediction and the last line is the experimental result.
[…]
Table 3 presents the lattice predictions for the RGTMEs of 6He $\beta$-decay, including the total and partial contributions from different orders of currents. Similar to 3H $\beta$-decay, the LO term remains dominant, while higher-order corrections contribute less than 5%. When compared with experiment, the LO result is slightly larger, and the negative correction from higher-order currents shifts the RGTME towards the correct direction and improves its agreement with the experimental data.
Since the contribution of higher-order currents are only included with the zeroth-order perturbation theory, as presented in Eq. (33), it is necessary to assess the impact of this truncation of the perturbative series. Our calculations indicate that the first-order perturbative correction to the LO current remains below 10% of the zeroth-order contribution. Assuming this is also the case for higher-order currents, the truncation would affect the total RGTME by about 1%, which is under good control and doesn’t influence our main conclusions.
VI Summary
In this work, we investigate nuclear $\beta$-decays using lattice quantum Monte Carlo simulations within the framework of the nuclear lattice effective field theory (NLEFT). The three-body low-energy constants (LECs) $c_{D}$ and $c_{E}$ are determined from the binding energy and the half-life of 3H. Our prediction for 6He $\beta$-decay shows reasonable agreement with experimental data and validates the method employed in this work.
To conclude, this work establishes a foundation for investigating $\beta$-decays in light and medium-mass nuclei with $A\geq 4$ using NLEFT. We anticipate that the recently developed N3LO interaction Li et al. (2018); Elhatisari et al. (2024a); Liu et al. will improve the predictions. Additionally, our approach opens the door to studying other nuclear electroweak processes of physical interest Engel and Menéndez (2017); Cirigliano et al. (2018a, b); Hardy and Towner (2020b); Cirigliano et al. (2024a, b); Gennari et al. (2025) within a unified lattice field theory framework, complementing lattice QCD studies at the quark and gluon level Tuo et al. (2019); Detmold and Murphy (2020); Davoudi et al. (2024); Feng et al. (2020); Ma et al. (2024b).
Acknowledgements.
We are grateful for helpful discussions on the nuclear interactions with Dean Lee, Ulf-G. Meißner and Shihang Shen. We appreciate Hermann Krebs for the choice of axial currents, Young-Ho Song for the design of Monte Carlo sampling method and Jiangming Yao for helpful comments. X.F. and T.W. were supported in part by NSFC of China under Grant No. 12125501, No. 12293060, No. 12293063, and No. 12141501, and National Key Research and Development Program of China under No. 2020YFA0406400. B.L. was supported by NSAF No.U2330401 and National Natural Science Foundation of China with Grant No.12275259.
Appendix A Nucleon-nucleon Scattering
We use the spherical wall method to calculate nucleon-nucleon phase shifts on the lattice SuppNN1 ; SuppNN2 ; SuppNN3 . The fitting result of $n$-$p$ scattering for the full Hamiltonian $H$ (blue dash-dot line) and the non-perturbative Hamiltonian $H_{0}$ (black dashed line) is shown in Fig. S 1, with the center-of-mass momentum $P\leq 200$ MeV. We also show the phase shifts extracted from the Nijmegen partial wave analysis SuppNN4 represented by the red points. For $H_{0}$, it is only fitted to ${}^{1}S_{0}$ and ${}^{3}S_{1}$ channels, so reasonable agreement with the experimental data appears in these two channels. The results for other channels are predictions based on the $S$-waves fit, and the agreement is not very satisfactory. We could in principle make an improvement by increasing the value $\Lambda^{\prime}_{\pi}$ in Eq. (13) or adding more operators in $H_{0}$, but this would introduce additional sign problems. For $H$, it shows good agreement with the experiment in most channels. In several $D$-wave channels there are some deviations, which can be eliminated if operators beyond N2LO are included.
Figure S 1: $n$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ for $H$ and $H_{0}$. The red points are from Nijmegen partial wave analysis. The blue and black lines denote fitting results of the full Hamiltonian $H$ and the non-perturbative Hamiltonian $H_{0}$, respectively.
Figure S 2: $p$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ in ${}^{1}S_{0}$ channel. The black points with errors are from Nijmegen partial wave analysis. The gray line denotes the fit result.
For the isospin-breaking potential $V_{pp}$ and $V_{nn}$, we tune the value of $C_{pp}$ to reproduce the $p$-$p$ phase shifts in ${}^{1}S_{0}$ channel SuppNN4 . The fitting result is given in Fig. S 2. For $V_{nn}$, we fit it to the experimental value of the $n$-$n$ scattering length SuppNN6 and effective range SuppNN7 in ${}^{1}S_{0}$ channel:
$a_{nn}=-18.95\pm 0.40\ \mathrm{fm},\quad r_{nn}=2.75\pm 0.11\ \mathrm{fm}.$ (40)
Appendix B Perturbative Expression of the RGTME
In this section, we give detailed expressions for the perturbative expansion of the RGTME.
We write Eq. (29) more explicitly as,
$\mathrm{GT}(t)=\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi_{f}|M^{N_{t}/2}\mathcal{J}M^{N_{t}/2}|\Phi_{i}\rangle}{\sqrt{\langle\Phi_{i}|M^{N_{t}}|\Phi_{i}\rangle}\sqrt{\langle\Phi_{f}|M^{N_{t}}|\Phi_{f}\rangle}}.$ (41)
According to Eq. (30), the full transfer matrix is separated into two parts,
$M=M_{0}+\Delta M.$ (42)
Then Eq. (41) can be expanded as a power series of $\Delta M$:
$\mathrm{GT}(t)=\mathrm{GT}^{(0)}(t)+\mathrm{GT}^{(1)}(t)+\mathcal{O}(\Delta M^{2}).$ (43)
To write down the expressions for $\mathrm{GT}^{(0)}$ and $\mathrm{GT}^{(1)}$, we introduce the notation
$|\Phi_{i/f,0}^{t/2}\rangle=M_{0}^{N_{t}/2}|\Phi_{i/f}\rangle$ (44)
to describe the evolution of trial states under the non-perturbative Hamiltonian $H_{0}$. We also introduce
$|\delta\Phi_{i/f}^{t/2}\rangle=\sum_{k=1}^{N_{t}/2}M_{0}^{k-1}\Delta MM_{0}^{N_{t}/2-k}|\Phi_{i/f}\rangle$ (45)
as the first-order correction to Eq. (44). With Eq. (44) and (45), it easy to derive,
$\mathrm{GT}^{(0)}(t)=\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi_{f,0}^{t/2}|\mathcal{J}|\Phi_{i,0}^{t/2}\rangle}{\sqrt{\langle\Phi_{i,0}^{t/2}|\Phi_{i,0}^{t/2}\rangle\langle\Phi_{f,0}^{t/2}|\Phi_{f,0}^{t/2}\rangle}}$ (46)
and
$\mathrm{GT}^{(1)}(t)$ $=$ $\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi_{f,0}^{t/2}|\mathcal{J}|\delta\Phi_{i}^{t/2}\rangle+\langle\delta\Phi_{f}^{t/2}|\mathcal{J}|\Phi_{i,0}^{t/2}\rangle}{\sqrt{\langle\Phi_{i,0}^{t/2}|\Phi_{i,0}^{t/2}\rangle\langle\Phi_{f,0}^{t/2}|\Phi_{f,0}^{t/2}\rangle}}$ (47)
-- $\mathrm{GT}^{(0)}(t)\left[\frac{\langle\Phi_{i,0}^{t/2}|\delta\Phi_{i}^{t/2}\rangle+\langle\delta\Phi_{i}^{t/2}|\Phi_{i,0}^{t/2}\rangle}{2\langle\Phi_{i,0}^{t/2}|\Phi_{i,0}^{t/2}\rangle}+i\leftrightarrow f\right].$
Appendix C More about the Monte Carlo Method
In this work, we need to calculate matrix elements of the following form,
$\langle\Psi^{t/2}_{a,0}|\mathcal{O}|\Psi^{t/2}_{b,0}\rangle,\ a,b\in\{i,f\},\ \mathcal{O}\in\{1,\mathcal{J}\},$ (48)
as is shown in Eq. (46) and (47). Applying auxiliary field transformation, we have,
$\langle\Phi_{a,0}^{t/2}|\mathcal{O}|\Phi_{b,0}^{t/2}\rangle=\int\mathcal{D}s\ e^{-S[s]}Z_{a,b}^{\mathcal{O}}[s].$ (49)
Here $s$ is the auxiliary field located at all the lattice points and $S[s]$ is a Gaussian-type function. The interaction between the nucleons and the auxiliary field is included in $Z_{a,b}^{O}[s]$. After the transformation, both the nominators and the denominators in Eq. (46),(47) become integrals over $s$. We evaluate them through generating a set of auxiliary fields which distribute according to a given probability distribution function $P[s]$. A proper choice of $P[s]$ can help improve the sampling efficiency and reduce statistical noises. As both the initial and the final state nucleus contribute to the $\beta$-decay process, we take the following form of $P[s]$:
$P[s]\propto c_{i}e^{-S[s]}|Z^{\mathcal{O}=1}_{i,i}[s]|+c_{f}e^{-S[s]}|Z^{\mathcal{O}=1}_{f,f}[s]|,$ (50)
with $c_{i}$ and $c_{f}$ positive numbers. Their values are tuned to suppress statistical errors. The update of the auxiliary field is accelerated by the shuttle algorithm proposed in Ref. Supp2 .