### v2

# Investigating nuclear $\beta$ decay using a lattice quantum Monte Carlo approach

Abstract
We present an ab initio calculation of nuclear $\beta$ decay within the framework of nuclear lattice effective field theory (NLEFT), employing auxiliary-field quantum Monte Carlo methods to solve the nuclear many-body problem. Our approach combines next-to-next-to-leading order two- and three-body chiral interactions with one- and two-body axial current operators, all consistently derived in chiral effective field theory. Low-energy constants are determined exclusively from nucleon-nucleon scattering phase shifts and few-body observables for systems with $A\leq 3$. Using these interactions and transition operators, we perform two-channel Monte Carlo simulations to compute the $\beta$-decay matrix element for 6He, obtaining results in reasonable agreement with experimental measurements. To address the Monte Carlo sign problem, we implement a perturbative expansion around a leading-order Hamiltonian with approximate Wigner-SU(4) symmetry. This systematic approach provides a foundation for extending NLEFT simulations to precision studies of weak processes in medium-mass nuclei.
I Introduction
Nuclear $\beta$ decay is a fundamental process crucial to understanding the laws of nature. Historically, studies of $\beta$ decay have led to groundbreaking discoveries, including the existence of neutrinos Brown (1978), the confirmation of parity violation Wu et al. (1957), and contributions to the unification of electroweak interactions Glashow (1961); Weinberg (1967); Salam (1968). Today, nuclear $\beta$ decay provides a sensitive probe for physics beyond the standard model. For a comprehensive review of ongoing and planned experimental programs, see Ref. González-Alonso et al. (2019) and references therein.
Besides its central role in particle physics, nuclear $\beta$ decay is also of great interest within the field of low-energy nuclear physics. Currently there remain several unresolved challenges in precisely describing the nuclear weak decay with the nucleonic degrees of freedom. For instance, traditional shell-model calculations tend to overestimate the reduced Gamow-Teller matrix element (RGTME), especially for medium-mass nuclei Brown and Wildenthal (1985); Martínez-Pinedo et al. (1996); Langanke et al. (1995). To reconcile this discrepancy, the axial charge $g_{A}$ in the axial current is often tuned to match experimental data Chou et al. (1993), leading to the well-known $g_{A}$ quenching problem. Moreover, the nuclear force binds nucleons into atomic nuclei, inducing complicated many-body correlations, which requires tremendous amount of computing resources and advanced algorithms to solve the corresponding quantum many-body problems. Lastly, the external electroweak field interacts with the nucleus via not only one-body currents but also many-body currents, so a proper description of $\beta$ decays requires a consistent treatment of the electroweak currents together with the complicated nuclear chiral interactions.
Recent years have seen rapid development of abinitioab\ initio nuclear many-body methods Carlson et al. (2015); Barrett et al. (2013); Hagen et al. (2014); Hergert et al. (2016); Lee (2009), which provide a consistent and systematically improvable framework for understanding the intricate interplays among nuclear forces, electroweak currents, and quantum many-body correlations in $\beta$-decay processes. For nuclei of $s$ and $p$ shells, calculations based on the quantum Monte Carlo (QMC) method Pastore et al. (2018); King et al. (2020) and the no-core shell model method Navrátil et al. (2007); Gysbers et al. (2019) have been performed systematically, with the latter applied to the study of the anomalous long lifetime of 14Maris et al. (2011). For heavier nuclei, the coupled-cluster method has been applied to investigating the impact of three-body forces and two-body currents on the Gamow-Teller transitions of 14C, 22O and 24Ekström et al. (2014), and it also provided the first abab initoinito calculation of 100Sn $\beta$ decay Gysbers et al. (2019). Besides, the valence-space in-medium similarity renormalization group method has made significant progress in studying $\beta$ decays of medium-mass nuclei Gysbers et al. (2019); Stroberg (2021); Yuan et al. (2025). These abab initioinitio calculations have significantly improved our understanding of nuclear $\beta$ decays, yet also raised a lot of open questions. For instance, the QMC calculations struggle to reproduce the experimental $\beta$ decay rate for 8Li, 8B, and 8He King et al. (2020). Furthermore, the contribution of two-body currents is found to have opposite signs in Refs. King et al. (2020) and Gysbers et al. (2019). This theoretical inconsistency was further explored in Ref. Gnech et al. (2021), which also revealed that the one-pion-exchange (OPE) current is highly sensitive to the nuclear forces and regularization schemes.
As one of the advanced nuclear ab initio methods, the nuclear lattice effective field theory (NLEFT) Lee (2009); Lähde and Meißner (2019) implements the chiral interactions on a cubic spatial lattice and solves the corresponding discrete many-body Schrödinger equation using the auxiliary-field Monte Carlo method. Recently the NLEFT has achieved remarkable progress in studying the properties of nuclear ground states and excited states Borasoy et al. (2007); Epelbaum et al. (2010a, b); Lähde et al. (2014); Lu et al. (2019, 2022); Elhatisari et al. (2024a); Epelbaum et al. (2014); Shen et al. (2023); Meißner et al. (2024); Shen et al. (2025); Zhang et al. ; Wang et al. , nuclear clustering Epelbaum et al. (2011, 2012, 2013); Elhatisari et al. (2016a, 2017), scattering processes Bour et al. (2012); Elhatisari et al. (2015), thermal nuclear matter Lu et al. (2020); Ren et al. (2024); Ma et al. (2024a) , and hypernuclei Bour et al. (2015); Frame et al. (2020); Hildenbrand et al. (2024). In this area, in recent years a significant amount of effort was made to build high-precision lattice interactions and design algorithms that mitigate the Monte Carlo sign problem. For example, the wave function matching method Elhatisari et al. (2024a) and perturbative quantum Monte Carlo method Lu et al. (2022); Liu et al. (2025) were proposed to improve the accuracy of perturbative calculations around a simple Hamiltonian without the sign problem. When combined with high-fidelity nuclear forces, these methods were able to reproduce the experimental binding energies and charge radii with systematically controllable statistical errors.
Attempts to study nuclear $\beta$ decay through NLEFT have just begun very recently. The first and currently the only study focuses on the triton $\beta$ decay Elhatisari et al. (2024b), which employs the recently developed high-fidelity next-to-next-to-next-to-leading order (N3LO) chiral interaction Elhatisari et al. (2024a) and the leading-order (LO) axial current to achieve consistent predictions for the Fermi and Gamow-Teller matrix elements. The sensitivity of these matrix elements to the strengths of three-body forces is also carefully examined therein.
In this work, we extend the work of Elhatisari et al. (2024b) by presenting an in-depth NLEFT study of nuclear $\beta$ decay, with major developments of two aspects. First, we have incorporated one- and two-body axial currents beyond LO in the calculation, which are essential for the determination of three-body low-energy constants (LECs), $c_{D}$ and $c_{E}$, of the next-to-next-to-leading order (N2LO) chiral nuclear force through triton $\beta$ decay  Gazit et al. (2009), and it has been shown that the two-body currents provide sizable contributions to Gamow-Teller matrix elements of certain nuclei King et al. (2020); Gysbers et al. (2019). Second, we have succeeded in simulating nuclear $\beta$ decay through the lattice quantum Monte Carlo method, which can be directly applied to weak decays of medium-mass and heavy nuclei. As a first step, here we focus on the two simplest $\beta$ decay processes, i.e., 3H $\rightarrow{{}^{3}}$He and 6He $\rightarrow{{}^{6}}$Li. We first fix $c_{D}$ and $c_{E}$ through the former process, then validate the applicability of our method using the latter. Since the main goal of this work is to build the relevant Monte Carlo machinery for investigating nuclear $\beta$ decay through NLEFT, we use the N2LO lattice chiral interaction fitted to nucleon-nucleon scattering phase shifts and few-body observables with $A\leq 3$ for numerical convenience. A systematic study based on the high-fidelity N3LO interaction is left for the future.
This paper is organized as follows. The nuclear interaction and axial currents used in this work are described in Secs. II and III. Our calculation methods and techniques are briefly discussed in Sec. IV. Results for 3H and 6He $\beta$ decay are presented in Sec. V. More details are provided in the Appendixes.
II Interaction
II.1 Full interaction
The full Hamiltonian used in this work is the N2LO chiral Hamiltonian discretized on a lattice with a lattice spacing $a=1.64$ fm,
$H=K+V_{\mathrm{OPE}}^{\Lambda_{\pi}}+V_{C_{\pi}}^{\Lambda_{\pi}}+V_{2\mathrm{N}}^{Q^{2}}+V_{\mathrm{IB}}+V^{Q^{3}}_{3\mathrm{N}}.$ (1)
Here, $K$ is kinetic term, and the nucleon mass is taken as $M_{N}=938.92$ MeV. The OPE potential is locally regularized in momentum space, following the prescription of Ref. Reinert et al. (2018),
$V_{\mathrm{OPE}}^{\Lambda_{\pi}}=-\frac{\tilde{g}_{A}^{2}}{4f_{\pi}^{2}}(\bm{\tau}_{1}\cdot\bm{\tau}_{2})\frac{\bm{(\sigma}_{1}\cdot\bm{q})(\bm{\sigma}_{2}\cdot\bm{q})}{\bm{q}^{2}+M_{\pi}^{2}}e^{-\frac{\bm{q}^{2}+M_{\pi}^{2}}{\Lambda_{\pi}^{2}}},$ (2)
$V_{C_{\pi}}^{\Lambda_{\pi}}=-\frac{\tilde{g}_{A}^{2}C_{\pi}}{4f_{\pi}^{2}}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})(\bm{\tau}_{1}\cdot\bm{\tau}_{2}),$ (3)
where
$C_{\pi}=\frac{\Lambda_{\pi}(\Lambda_{\pi}^{2}-2M_{\pi}^{2})+2\sqrt{\pi}M_{\pi}^{3}\exp\left(\frac{M_{\pi}^{2}}{\Lambda_{\pi}^{2}}\right)\mathrm{erfc}\left(\frac{M_{\pi}}{\Lambda_{\pi}}\right)}{3\Lambda_{\pi}^{3}}.$ (4)
The numerical values of the parameters used are $\tilde{g}_{A}=1.287$ for the nucleon axial charge, $f_{\pi}=92.2$ MeV for the pion decay constant, $M_{\pi}=134.98$ MeV for the pion mass, and $\Lambda_{\pi}=300$ MeV for the local regulator. The axial charge $\tilde{g}_{A}$ incorporates the Goldberger-Treiman discrepancy Arndt et al. (1994) and differs slightly from value $g_{A}=1.2723(23)$ Tanabashi et al. (2018) in the axial current. To distinguish between them, we use a tilde notation. The term in Eq. (3) serves as a counterterm to remove the short-distance singularity in the OPE potential. It is worth noting that the characteristic momentum scale in this work is around 100 MeV, considerably lower than twice the pion mass. As a result, the two-pion exchange (TPE) potential, which first appears at next-to leading order (NLO), cannot be distinguished from contact terms Epelbaum et al. (2009).
The two-nucleon (2N) short-range potential, $V^{Q^{2}}_{2\mathrm{N}}$, includes two contact terms at leading order (LO) and seven additional contact terms at next-to-leading order (NLO),
$V_{2\mathrm{N}}^{Q^{2}}$ $=$ $B_{1}+B_{2}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})+C_{1}q^{2}+C_{2}q^{2}(\bm{\tau}_{1}\cdot\bm{\tau}_{2})$ (5)
$+$ $C_{3}q^{2}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})+C_{4}q^{2}(\bm{\sigma}_{1}\cdot\bm{\sigma}_{2})(\bm{\tau}_{1}\cdot\bm{\tau}_{2})$
$+$ $C_{5}(\bm{\sigma}_{1}\cdot\bm{q})(\bm{\sigma}_{2}\cdot\bm{q})+C_{6}(\bm{\sigma}_{1}\cdot\bm{q})(\bm{\sigma}_{2}\cdot\bm{q})(\bm{\tau}_{1}\cdot\bm{\tau}_{2})$
$+$ $\frac{i}{2}C_{7}(\bm{q}\times\bm{k})\cdot(\bm{\sigma}_{1}+\bm{\sigma}_{2}),$
where $B_{i}$ and $C_{i}$ are LECs fitted to the Nijmegen neutron-proton ($n$-$p$) scattering phase shifts Stoks et al. (1993). All contact terms are regularized using a nonlocal Gaussian regulator in momentum space, following Ref. Lu and Deng . Specifically, for the matrix element $V(\bm{p}_{1},\bm{p}_{1}^{\prime},\bm{p}_{2},\bm{p}_{2}^{\prime})$, where $\bm{p}_{i}$ and $\bm{p}^{\prime}_{i}$ are the incoming and outgoing momenta of the $i$th nucleon, we apply a non-local momentum regulator
$f_{2\mathrm{N}}=\prod_{i=1}^{2}\exp\left(-\frac{p_{i}^{6}+p_{i}^{\prime 6}}{2\Lambda^{6}}\right).$ (6)
In the center-of-mass frame, Eq. (6) reduces to the nonlocal regulator commonly used in Ref.  Entem and Machleidt (2003).
The isospin-breaking term, $V_{\mathrm{IB}}$, first appears at NLO. Following the convention of Ref. Epelbaum et al. (2010a), we separate it into three components: the long-range Coulomb term,
$V_{\mathrm{IB}}^{\mathrm{Cou}}=\frac{\alpha_{\mathrm{EM}}}{\mathrm{max}(|\bm{r}_{1}-\bm{r}_{2}|,0.5)}\left(\frac{1+\tau_{1,z}}{2}\right)\left(\frac{1+\tau_{2,z}}{2}\right),$ (7)
the short-range proton-proton ($p$-$p$) contact term,
$V_{pp}=C_{pp}\delta(\bm{r}_{1}-\bm{r}_{2})\left(\frac{1+\tau_{1,z}}{2}\right)\left(\frac{1+\tau_{2,z}}{2}\right),$ (8)
and the short-range neutron-neutron ($n$-$n$) contact term,
$V_{nn}=C_{nn}\delta(\bm{r}_{1}-\bm{r}_{2})\left(\frac{1-\tau_{1,z}}{2}\right)\left(\frac{1-\tau_{2,z}}{2}\right).$ (9)
These isospin-breaking terms are essential for reproducing the 3H-3He mass splitting Epelbaum et al. (2010b). The 2N LECs $C_{pp}$ and $C_{nn}$ in Eqs. (8) and (9) are fitted to experimental $p$-$p$ phase shifts, as well as $n$-$n$ scattering lengths and effective ranges Gonzalez Trotter et al. (2006); Miller et al. (1990). Details on determining the LECs are provided in Appendix A.
Finally, for the three-nucleon (3N) potential $V_{3\mathrm{N}}^{Q^{3}}$, which first appears at N2LO, we include the OPE term
$V_{3\mathrm{N}}^{\mathrm{OPE}}$ $=$ $-\frac{g_{A}}{8f^{2}_{\pi}}\frac{c_{D}}{f_{\pi}^{2}\Lambda_{\chi}}\sum_{C(i,j,k)}\frac{(\bm{\sigma}_{i}\cdot\bm{q}_{i})(\bm{\sigma}_{j}\cdot\bm{q}_{i})}{\bm{q}_{i}^{2}+M_{\pi}^{2}}(\bm{\tau}_{i}\cdot\bm{\tau}_{j}),$
and the 3N contact term
$V_{3\mathrm{N}}^{\mathrm{CT}}=\frac{c_{E}}{2f_{\pi}^{4}\Lambda_{\chi}}\sum_{C(i,j,k)}\bm{\tau}_{i}\cdot\bm{\tau}_{j}.$ (11)
Here, $\Lambda_{\chi}=700$ MeV is the chiral symmetry breaking scale. The 3N regulator
$f_{3\mathrm{N}}=\prod_{i=1}^{3}\exp\left(-\frac{p_{i}^{6}+p_{i}^{\prime 6}}{2\Lambda^{6}}\right)$ (12)
is employed to regularize $V_{3\mathrm{N}}^{\mathrm{CT}}$. We have also tested the 3N TPE potential but found its contribution negligible, so we neglect it in this work. The LECs $c_{D}$ and $c_{E}$ are determined by fitting to the experimental 3H binding energy and the RGTME of 3H $\beta$-decay process simultaneously. Details are given in Sec. V.
II.2 Nonperturbative interaction
The full Hamiltonian in Eq. (1) includes complex spin-isospin-dependent terms, which pose significant challenges for lattice calculations. One major issue is that these terms can induce severe sign problems in fermionic systems, such as nucleons, in Monte Carlo simulations Lähde and Meißner (2019). To mitigate this, one can construct a simplified Hamiltonian $H_{0}$ which captures the essential physics but with much milder sign problems, and solve the full Hamiltonian $H$ perturbatively Lu et al. (2022). Details of the perturbative method will be provided in Sec. IV.1. In this work, we consider the following nonperturbative Hamiltonian:
$H_{0}$ $=$ $K+\frac{C_{2}}{2!}\sum_{\bm{n}}:\tilde{\rho}^{2}(\bm{n}):+\frac{C_{3}}{3!}\sum_{\bm{n}}:\tilde{\rho}^{3}(\bm{n}):$ (13)
$+$ $\frac{C_{I}}{2!}\sum_{I,\bm{n}}:\tilde{\rho}^{2}_{I}(\bm{n}):+V_{\mathrm{OPE}}^{\Lambda_{\pi}^{\prime}}.$
Here, :: represents normal ordering and $\tilde{\rho}$ is the SU(4)-symmetric density operator smeared both locally and nonlocally Elhatisari et al. (2017):
$\tilde{\rho}(\bm{n})$ $=$ $\sum_{s,i}\tilde{a}^{\dagger}_{s,i}(\bm{n})\tilde{a}_{s,i}(\bm{n})$ (14)
$+$ $s_{\mathrm{L}}\sum_{|\bm{n}-\bm{n}^{\prime}|=1}\sum_{s,i}\tilde{a}^{\dagger}_{s,i}(\bm{n}^{\prime})\tilde{a}_{s,i}(\bm{n}^{\prime}),$
where $s$ and $i$ are the spin and isospin indexes, respectively. The parameter $s_{\mathrm{L}}$ controls the range of the local part of the interaction. The smeared creation operator is defined as
$\tilde{a}^{\dagger}_{s,i}(\bm{n})=a^{\dagger}_{s,i}(\bm{n})+s_{\mathrm{NL}}\sum_{|\bm{n}-\bm{n}^{\prime}|=1}\tilde{a}^{\dagger}_{s,i}(\bm{n}^{\prime}),$ (15)
with $s_{\mathrm{NL}}$ controlling the range of the nonlocal part of the interaction. The definition of the isospin-dependent density operator $\tilde{\rho}_{I}$ is similar to $\tilde{\rho}$:
$\tilde{\rho}_{I}(\bm{n})$ $=$ $\sum_{s}\sum_{i,i^{\prime}}\tilde{a}^{\dagger}_{s,i}(\bm{n})[\tau_{I}]_{i,i^{\prime}}\tilde{a}_{s,i^{\prime}}(\bm{n})$
$+$ $s_{\mathrm{L}}\sum_{|\bm{n}-\bm{n}^{\prime}|=1}\sum_{s}\sum_{i,i^{\prime}}\tilde{a}^{\dagger}_{s^{\prime},i}(\bm{n}^{\prime})[\tau_{I}]_{i,i^{\prime}}\tilde{a}_{s^{\prime},i^{\prime}}(\bm{n}^{\prime}),$
with $\tau_{I}$ the $I$th Pauli matrix .
The Hamiltonian in Eq. (13) consists of two parts, written separately for clarity. The first line contains terms that respect Wigner’s SU(4) symmetry Wigner (1937), an approximate symmetry of low-energy nuclear interactions. The inclusion of this symmetry helps suppress sign problems in lattice Monte Carlo calculations Chen et al. (2004). Recently, the SU(4)-symmetric Hamiltonian has been applied to the study of the properties of light and medium-mass nuclei as well as the equation of state of symmetric nuclear matter and neutron matter Lu et al. (2019), the intrinsic clustering geometry of 12C’s low-lying states Shen et al. (2023), the thermodynamics of nuclear systems Lu et al. (2020), the clustering effects in hot, dilute nuclear matter Ren et al. (2024), and the $\alpha$-particle monopole transition form factor Meißner et al. (2024). The second line of Eq. (13) introduces an isospin-dependent contact term and a OPE term, both of which explicitly break the SU(4) symmetry. These additional terms are introduced to improve the overall description of $p$-shell nuclei Wiringa and Pieper (2002) , such as 6He and 6Li studied in this work. To suppress the sign problem caused by the tensor force, we adopt a lower cutoff of $\Lambda_{\pi}^{\prime}=180$ MeV for the OPE term in Eq. (13), compared to the higher cutoff of $\Lambda_{\pi}=300$ MeV used in the full Hamiltonian in Eq. (1).
To determine the unknown parameters in Eq. (13), we perform a combined fit to $n$-$p$ scattering phase shifts in two $S$-wave channels, along with the ground-state binding energies of 4He and 6He. The inclusion of the latter two observables is motivated by the fact that 6He is a weakly bound halo nucleus Tanihata (1996); Jensen et al. (2004); Zhukov et al. (1993), with its binding energy very close to that of 4He:
$E(^{4}\mathrm{He})=-28.30\mathrm{MeV},\ E(^{6}\mathrm{He})=-29.27\mathrm{MeV}.$ (17)
Therefore, we carefully tune $H_{0}$ to ensure that 6He is bound in our calculation. The fitted parameter values are presented in Table 1.
$s_{\mathrm{L}}$ $s_{\mathrm{NL}}$ $C_{2}(\mathrm{MeV}^{-2})$ $C_{I}(\mathrm{MeV}^{-2})$ $C_{3}(\mathrm{MeV}^{-5})$
[…]
Table 1: Values for all the parameters in Eq. (13).
III Axial Current
The transition operator for the RGTME is related to the nuclear axial current. We employ one- and two-body $\Delta$-less chiral EFT axial currents derived from the unitary transformation method of the Bonn group Krebs et al. (2017) and the time-ordered perturbation method of the JLab-Pisa group Baroni et al. (2016a) . These two approaches yield consistent tree-level results up to N3LO ($Q^{0}$) but exhibit some differences at N4LO ($Q$) due to loop contributions Baroni et al. (2016b); Krebs et al. (2017, 2020). Moreover, Ref. Krebs (2020) recently highlighted that the absence of consistent regularization for axial-vector current operators leads to a violation of chiral symmetry in the chiral limit at N4LO. For these reasons, we restrict our analysis to tree-level axial currents up to N3LO and neglect higher-order contributions.
Figure 1: Diagrams illustrating the contributions to the axial current up to N3LO used in this work. Nucleons, pions and the external field are denoted by the solid , dashed, and wavy lines respectively. Panel (a) denotes the LO current while panel (b) is the relativistic correction at N2LO. Panels (c) and (d) are the OPE current and the contact current at N3LO, respectively.
Since the momentum $\bm{k}$ carried by the current is very small in $\beta$-decay processes, we work in the $\bm{k}=0$ limit to simplify calculations. Under this limit, there are in total four terms that contribute to the axial current up to N3LO: the one-body current $\bm{A}_{\mathrm{LO}}$ at LO , the relativistic correction term $\bm{A}_{\mathrm{N}^{2}\mathrm{LO}}$ at N2LO, the OPE term $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$, and the contact term $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})$ at N3LO. Figure 1 provides a diagrammatic illustration of them, and their expressions in the momentum space can be found in Refs. Baroni et al. (2016b); Park et al. (2003). The OPE current involves two nucleon-pion coupling constants, $c_{3}$ and $c_{4}$, for which we take the values $c_{3}=-5.61$ GeV-1 and $c_{4}=4.26$ GeV-1 Hoferichter et al. (2015). The contact current is related to the three-body force coefficient $c_{D}$ as follows Park et al. (2003); Baroni et al. (2018):
$\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})=-\frac{c_{D}}{8f_{\pi}^{2}\Lambda_{\chi}}\sum_{i<j}^{A}(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}(\bm{\sigma}_{i}\times\bm{\sigma}_{j}),$ (18)
where we used the shorthand notation $(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}=(\bm{\tau}_{i}\times\bm{\tau}_{j})^{x}+i(\bm{\tau}_{i}\times\bm{\tau}_{j})^{y}$ for the two-body isospin-raising operator.
The realization of the axial current on the lattice is straightforward. Given the momentum-space expression of the current, the finite lattice spacing and box size provides natural ultraviolet and infrared cutoffs, respectively, which are used to discretize the momenta and constrain them in the first Brillouin zone. The corresponding operator in the configuration space is then constructed through fast Fourier transformation.
As a comparison, in the following we briefly discuss the treatment of two-body currents for calculations in the continuum Park et al. (2003); Gazit et al. (2009); Baroni et al. (2018); King et al. (2020); Gnech et al. (2021). These calculations rely on analytical configuration-space expressions of the OPE and contact currents, which look like
$\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$ $=$ $\frac{g_{A}}{2f_{\pi}^{2}}\left(\frac{1}{3}c_{3}+\frac{2}{3}c_{4}+\frac{1}{6M_{N}}\right)\sum_{i<j}^{A}\delta^{(3)}(\bm{r}_{ij})$
$\times(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}(\bm{\sigma}_{i}\times\bm{\sigma}_{j})+\tilde{\bm{A}}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$
$\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})$ $=$ $-\frac{c_{D}}{8f_{\pi}^{2}\Lambda_{\chi}}\sum_{i<j}^{A}\delta^{(3)}(\bm{r}_{ij})(\bm{\tau}_{i}\times\bm{\tau}_{j})^{+}(\bm{\sigma}_{i}\times\bm{\sigma}_{j}).$
The above equations can be easily derived by performing Fourier transformation on momentum-space expressions of the current. Note that $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{OPE})$ consists of two parts: the first part proportional to $\delta^{(3)}(\bm{r}_{ij})=\delta^{(3)}(\bm{r}_{i}-\bm{r}_{j})$ absorbs the short-range singularity, while the second part with a tilde $\tilde{\bm{A}}_{\mathrm{N}^{3}\mathrm{LO}}$(OPE) collects remaining long-range contributions, whose explicit form can be found in Refs. Pastore et al. (2018); King et al. (2020); Gnech et al. (2021). Because of the Fermi-Dirac statistics Park et al. (2003), the first part has the same spin-isopsin structure as the contact current $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}(\mathrm{CT})$. Therefore, some works absorb the former into the latter and call the remaining term $\tilde{\bm{A}}_{\mathrm{N}^{3}\mathrm{LO}}$(OPE) as the OPE current Park et al. (2003); Gazit et al. (2009); Baroni et al. (2018); King et al. (2020); Gnech et al. (2021). There is also literature which defines the full operator $\bm{A}_{\mathrm{N}^{3}\mathrm{LO}}$(OPE) as the OPE current Baroni et al. (2016b). We mention that these two definitions are equivalent to each other, and would lead to consistent predictions on physical observables such as the half-life. For our lattice calculation, we find it more convenient to follow the second prescription, which makes it easier to transform the current between the momentum representation and configuration representation, numerically.
IV Computational Methods
In this section, we introduce two methods used in this work, i.e., the Monte Carlo method and the sparse-matrix diagonalization method, to calculate the RGTME following the convention of Edmonds Edmonds (1960),
$\mathrm{GT}=\frac{\sqrt{2J_{f}+1}}{g_{A}}\frac{\langle J_{f}M_{f}|\mathcal{J}|J_{i}M_{i}\rangle}{\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}.$ (20)
Here, $|J_{i}M_{i}\rangle$ and $|J_{f}M_{f}\rangle$ represent the eigenstates of the full Hamiltonian $H$ for the initial and final nucleus, respectively, with $J$ the nuclear spin and $M$ the magnetic quantum number. The weak transition operator $\mathcal{J}$ is defined as
$\mathcal{J}=-\frac{1}{\sqrt{2}}(A_{x}+iA_{y}),$ (21)
[…]
We comment that the definition of the RGTME, Eq. (20), is followed throughout the work. The same convention has also been adopted to study 6He $\beta$ decay in Ref. Gnech et al. (2021); King et al. (2020), so our result of 6He’s RGTME can be directly compared to theirs. For 3H $\beta$ decay, however, different conventions have been adopted by different works. For example, our definition of the RGTME is larger than the Gamow-Teller matrix element defined in Ref. Baroni et al. (2016b, 2018); Gnech et al. (2021); Elhatisari et al. (2024b) by a factor of $\sqrt{2}$, but the same as the one defined in Ref. Li Muli et al. (2025). One should be aware of this difference when making a comparison.
For latter convenience, we decompose $\mathcal{J}$ into different orders according to Fig. 1,
$\mathcal{J}=\mathcal{J}_{\mathrm{LO}}+\mathcal{J}_{\mathrm{N}^{2}\mathrm{LO}}+\mathcal{J}_{\mathrm{N}^{3}\mathrm{LO}}.$ (22)
Accordingly, the RGTME can be expanded order by order as
$\mathrm{GT}=\mathrm{GT}_{\mathrm{LO}}+\mathrm{GT}_{\mathrm{N}^{2}\mathrm{LO}}+\mathrm{GT}_{\mathrm{N}^{3}\mathrm{LO}}.$ (23)
IV.1 Monte Carlo method
The Monte Carlo method is a crucial tool for simulating nuclear systems in NLEFT. In this section, we discuss its application to our work.
We begin by introducing the full transfer matrix
$M=:e^{-a_{t}H}:$ (24)
for the full Hamiltonian $H$ defined in Eq. (1), where $a_{t}$ is the temporal lattice spacing. To facilitate perturbative calculations that will be introduced below, we also define a nonperturbative transfer matrix
$M_{0}=:e^{-a_{t}H_{0}}:,$ (25)
corresponding to the nonperturbative Hamiltonian $H_{0}$ defined in Eq. (13).
Next, we prepare the trial states $|\Phi_{i/f}\rangle$, which share the same quantum numbers as $|J_{i/f}M_{i/f}\rangle$. These states are evolved along Euclidean time using the transfer matrix $M$,
$|\Phi^{t/2}_{i/f}\rangle\equiv M^{N_{t}/2}|\Phi_{i/f}\rangle,$ (26)
where $N_{t}$ is the number of evolution steps and $t=N_{t}a_{t}$ represents the evolution time. Since both processes studied here, i.e., $^{3}\mathrm{H}\rightarrow^{3}\mathrm{He}$ and $^{6}\mathrm{He}\rightarrow$ $^{6}\mathrm{Li}$, only involve ground-state nuclei, $t$ is taken as large as possible to suppress excited-state contamination. To accelerate the convergence, the trial states $|\Psi_{i/f}\rangle$ are constructed based on the the nuclear shell model Mayer (1948); Heyde (1994).
The RGTME, given in Eq. (20), can be extracted from the ratio
$\mathrm{GT}(t)=\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi^{t/2}_{f}|\mathcal{J}|\Phi^{t/2}_{i}\rangle}{\sqrt{\langle\Phi^{t/2}_{i}|\Phi^{t/2}_{i}\rangle\langle\Phi^{t/2}_{f}|\Phi^{t/2}_{f}\rangle}}$ (27)
at large $t$. To evaluate it, auxiliary field transformations are applied to both the numerator and denominator, reducing the problem to a multidimensional integral that can be handled using Monte Carlo methods. A detailed discussion of this process is provided in Appendix C. However, as discussed in Sec. II.2, a direct auxiliary field transformation on the full transfer matrix $M$ leads to severe sign problems, making the results highly noisy.
To address this issue, we employ perturbation theory Lu et al. (2022). Specifically, we decompose the transfer matrix into
$M=M_{0}+\Delta M,$ (28)
treating $\Delta M$ as a perturbation. This allows us to expand Eq. (27) as a perturbative series:
$\mathrm{GT}=\sum_{n=0}^{\infty}\mathrm{GT}^{(n)}_{\mathrm{LO}}+\sum_{n=0}^{\infty}\mathrm{GT}^{(n)}_{\mathrm{N}^{2}\mathrm{LO}}+\sum_{n=0}^{\infty}\mathrm{GT}^{(n)}_{\mathrm{N}^{3}\mathrm{LO}}$, (29)
with the time variable $t$ suppressed for brevity. Here, $n$ denotes the order of expansion in $\Delta M$. The contributions from different orders of the current are written as LO, N2LO, and N3LO, respectively. For the LO current, we retain the first two terms in the perturbative series and neglect higher-order corrections, namely,
$\mathrm{GT}_{\mathrm{LO}}\approx\mathrm{GT}^{(0)}_{\mathrm{LO}}+\mathrm{GT}^{(1)}_{\mathrm{LO}}$. (30)
For currents beyond LO, we keep only the first term,
$\mathrm{GT}_{\mathrm{N}^{2}\mathrm{LO}/\mathrm{N}^{3}\mathrm{LO}}\approx\mathrm{GT}^{(0)}_{\mathrm{N}^{2}\mathrm{LO}/\mathrm{N}^{3}\mathrm{LO}}$. (31)
Thus, Eq. (29) is approximated as
$\mathrm{GT}$ $\approx$ $\sum_{n=0}^{1}\mathrm{GT}^{(n)}_{\mathrm{LO}}+\mathrm{GT}^{(0)}_{\mathrm{N}^{2}\mathrm{LO}}+\mathrm{GT}^{(0)}_{\mathrm{N}^{3}\mathrm{LO}}$. (32)
The explicit expressions for the terms on the right-hand side of Eq. (32) are provided in Appendix B. They are computed using Monte Carlo methods. Since the system now evolves under $H_{0}$ that has an approximate Wigner-SU(4) symmetry instead of the full Hamiltonian $H$, the sign problem is significantly alleviated.
When designing the algorithm, we note that any $\beta$-decay process involves two different nuclei, making this essentially a two-channel problem. To control statistical errors, an appropriate probability distribution function is required to update the auxiliary fields during the Markov process. We leave relevant details to Appendix C.
IV.2 Lanczos eigenvector method
[…]
The main advantage of the Lanczos method is its accuracy. Unlike Monte Carlo simulations, it is free from statistical errors and systematic errors from the perturbative method. Additionally, due to its precision, this method enables a precise determination of the three-body LECs $c_{D}$ and $c_{E}$ from 3H $\beta$ decay. Specifically, we fit them to the experimental binding energy and half-life of 3H, closely following the procedure in Refs. Gazit et al. (2009); Baroni et al. (2016b, 2018); Elhatisari et al. (2024b). The details are presented in the next section.
V Results
V.1 3H $\beta$ decay
The details of 3H $\beta$ decay are discussed in this subsection. An essential step here is to convert the half-life, $t_{1/2}$, into the RGTME. We make use of the formula Simpson (1987); Hardy et al. (1990); Schiavilla et al. (1998); Chou et al. (1993)
$(1+\delta_{R})t_{1/2}f_{V}=\frac{2\pi^{3}\mathrm{ln}2/[m_{e}^{5}V_{ud}^{2}G_{F}^{2}(1+\Delta_{R}^{V})]}{B(\mathrm{F})+B(\mathrm{GT})f_{A}/f_{V}},$ (33)
which expresses $t_{1/2}$ in terms of the Fermi transition strength
$B(\mathrm{F})=\left(\sum_{n=1}^{3}\langle^{3}\mathrm{He}|\tau_{n,+}|^{3}\mathrm{H}\rangle\right)^{2},$ (34)
and the Gamow-Teller transition strength $B(\mathrm{GT})$, quadratic in the RGTME Chou et al. (1993),
$B(\mathrm{GT})=\frac{1}{2}g_{A}^{2}\mathrm{GT}^{2}.$ (35)
In Eq. (33), $m_{e}=0.511$ MeV is the electron mass, $V_{ud}=0.9737(3)$ is the up-down quark-mixing element of the Cabibbo-Kobayashi-Maskawa matrix Hardy and Towner (2020), $G_{F}=1.1664\times 10^{-5}$ GeV-1 is the Fermi constant Tishchenko et al. (2013), $f_{V}=2.8355\times 10^{-6}$ and $f_{A}=2.8505\times 10^{-6}$ are Fermi functions Simpson (1987), $\delta_{R}=1.9\%$ is the outer radiative correction of 3H $\beta$ decay Raman et al. (1978), and $\Delta_{R}^{V}=0.02467(22)$ is the inner radiative correction Seng et al. (2018). Our lattice result of the Fermi transition strength is
$B(\mathrm{F})=0.9996^{2}.$ (36)
Note that because of vector current conservation, $B(\mathrm{F})$ is strictly equal to 1 in the isospin-symmetric limit. In our calculation, the isospin-breaking potential $V_{\mathrm{IB}}$ introduced in Eqs. (7), (8) and (9) generates a small shift, at the level of $10^{-4}$. Combining the Fermi transition strength Eq. (36) with the measured comparative half-life, $(1+\delta_{R})t_{1/2}f_{V}=1134.6(31)$Simpson (1987), the value of the RGTME is determined as
$\mathrm{GT}_{\mathrm{exp}}=2.3297(58).$ (37)
The subscript $\mathrm{exp}$” means that the result is derived from experimental data, although the lattice prediction of the Fermi transition strength is used as an input. Note that our definition of 3H’s RGTME is different from Ref. Gazit et al. (2009); Baroni et al. (2016b, 2018); Elhatisari et al. (2024b) by an overall constant $\sqrt{2}$, as has been emphasized in the last section.
Now we show the determination of $c_{D}$ and $c_{E}$ using the Lanczos eigenvector method. In Fig. 2, $c_{E}$ is plotted against $c_{D}$ by fitting to the experimental binding energy $E(^{3}$H)=8.482)=-8.482 MeV. Similar to previous studies, we observe a nearly linear relation between $c_{D}$ and $c_{E}$ Gazit et al. (2009); Baroni et al. (2016b, 2018). Then we calculate the lattice RGTME, $\mathrm{GT}_{\mathrm{lat}}$, and plot its ratio with respect to $\mathrm{GT}_{\mathrm{exp}}$ in Fig. 3, as a function of $c_{D}$. All the points in Fig. 3 share the same values of $(c_{D},c_{E})$ with the ones in Fig. 2. Again, GTth/GTexp is a nearly linear function of $c_{D}$, and the values of $c_{D}$ and $c_{E}$ are determined as
$c_{D}=1.01(16),\quad c_{E}=0.81(1).$ (38)
Note that low-energy nucleon-deuteron ($n$-$d$) scattering provides an alternative process to help pin down $c_{D}$ and $c_{E}$ and to test the result above from 3H $\beta$ decay. The strategy has been employed by several works in the continuum  Epelbaum et al. (2002); Piarulli et al. (2018); Epelbaum et al. (2019), but it should be understood that a direct comparison of $c_{D}$ and $c_{E}$ cannot be performed due to different regularization schemes and regulators used in our work and in  Epelbaum et al. (2002); Piarulli et al. (2018); Epelbaum et al. (2019). In NLEFT, the adiabatic projection method  Rokash et al. (2015); Elhatisari (2019) was applied to investigating $n$-$d$ scattering Pine et al. (2013); Elhatisari et al. (2016b) using LO pionless EFT, with no discussion on the determination of three-body forces. Thus, a more systematic study of $n$-$d$ scattering on the lattice is required.
Figure 2: The $c_{E}$-$c_{D}$ trajectories obtained by fitting the experimental 3H binding energy 8.482 MeV.
Figure 3: The ratio GTlat/GTexp as a function of $c_{D}$. The gray band shows the experimental uncertainty of GTexp.
LO N2LO N3LO(OPE) N3LO(CT)
LE 2.3684 -0.0117 -0.0679 0.0409
MC 2.355(17) -0.010(2) -0.070(1) 0.039(3)
Table 2: Contributions to RGTME from currents at different orders for $c_{D}=$1.01, $c_{E}$=0.81. The upper and lower lines are the results of the Lanczos eigenvector method (LE) and the Monte Carlo (MC) method, respectively. The value in parentheses is the statistical error.
Nuclei($J^{P}$) 6He(0+) 6Li(1+)
EE/MeV $R_{\mathrm{ch}}$/fm EE/MeV $R_{\mathrm{ch}}$/fm
Lattice -28.80(1.21) 1.988(32) -31.88(93) 2.534(18)
Expt. -29.27 2.068(11) -31.99 2.589(39)
Table 3: Ground state energies and the charge radii of 6He and 6Li calculated on the lattice compared to experiment. The third line is the lattice prediction and the last line is the experimental result.
Figure 4: RGTME of 6He $\beta$-decay versus the Euclidean evolution time $t$. In the upper panel, the black points represent the contribution of the LO axial current, the red points denote the full RGTME $\mathrm{GT}$ and the red band is the fitting result. The experimental value of $\mathrm{GT}$ is plotted as the dashed purple line. The lower panel shows the net contributions from currents beyond LO.
LO N2LO N3LO(OPE) N3LO(CT) Total-LO Total Expt. Knecht et al. (2012)
RGTME 2.289(19) -0.024(2) -0.104(1) 0.032(1) -0.096(2) 2.192(20) 2.161(4)
Table 4: Contributions to RGTME of 6He $\beta$ decay from currents at different orders. Columns labeled with LO, N2LO, N3LO(OPE), and N3LO(CT) correspond to the diagrams illustrated in panels (a), (b), (c), and (d) of Fig. 1, respectively. The full result is given in the column labeled “Total,” while results including only corrections beyond LO are listed under “Total-LO.” The value in parenthese is the statistical uncertainty. The experimental value in the last column is from Ref. Knecht et al. (2012).
To evaluate the contributions from currents at different orders , we calculate the corresponding RGTMEs and present the results in Table 2. It is evident that the LO current provides the dominant contribution, while corrections from higher-order currents remain below 5%. To validate the two computational approaches described in Sec. IV, we also performed Monte Carlo calculations for 3H $\beta$ decay. The results of both approaches, shown in the second and third rows of Table 2, exhibit good agreement.
[…]
V.2 6He $\beta$ decay
We first present the energies and charge radii of 6He and 6Li in Table 3. The energy is computed up to second-order perturbation theory Lu et al. (2022), leading to relatively large uncertainties. The charge radius, on the other hand, is calculated up to the first order. Overall, our results show satisfactory agreement with experimental data.
Next, we present the results for 6He $\beta$ decay. The dependence of the RGTME on the Euclidean evolution time $t$ is illustrated in Fig. 4, where contributions from currents at LO (black points) and beyond LO (blue points) are plotted separately in the upper and lower panels, respectively. The full result is represented by the red points in the upper panel. For the LO current, we observe a monotonic increase with $t$ and find a plateau beginning at around $t=0.3$ MeV-1. In contrast, the net contribution beyond LO exhibits a relatively weaker dependence on $t$. Therefore, we make a constant fit in the interval 0.3 $\leq t\leq$ 0.5 MeV-1 for the LO results and 0.15 $\leq t\leq$ 0.5 MeV-1 for the beyond-LO results. The fit result for the full RGTME is plotted as the red band in the upper panel, and is very close to its experimental value represented by the horizontal dashed purple line.
Table 4 presents the lattice predictions for the RGTMEs of 6He $\beta$-decay, including the total and partial contributions from different orders of currents. Similarly to 3H $\beta$ decay, the LO term remains dominant, while higher-order corrections contribute less than 5%. When compared with experiment, the LO result is slightly larger, and the negative correction from higher-order currents shifts the RGTME towards the correct direction and improves its agreement with the experimental data.
Since the contributions of higher-order currents are only included with the zeroth-order perturbation theory, as presented in Eq. (31), it is necessary to assess the impact of this truncation of the perturbative series. Our calculations indicate that the first-order perturbative correction to the LO current remains below 10% of the zeroth-order contribution. Assuming this is also the case for higher-order currents, the truncation would affect the total RGTME by about 1%, which is under good control and does not influence our main conclusions.
VI Summary
In this work, we investigate nuclear $\beta$ decays using lattice quantum Monte Carlo simulations within the framework of the nuclear lattice effective field theory (NLEFT). The three-body low-energy constants (LECs) $c_{D}$ and $c_{E}$ are determined from the binding energy and the half-life of 3H. Our prediction for 6He $\beta$ decay shows reasonable agreement with experimental data and validates the method employed in this work.
To conclude, this work establishes a foundation for investigating $\beta$ decays in light and medium-mass nuclei with $A\geq 4$ using NLEFT. We anticipate that the recently developed N3LO interaction Li et al. (2018); Elhatisari et al. (2024a); Liu et al. (2025) will improve the predictions. Additionally, our approach opens the door to studying other nuclear electroweak processes of physical interest Engel and Menéndez (2017); Cirigliano et al. (2018a, b); Hardy and Towner (2020); Cirigliano et al. (2024a, b); Gennari et al. (2025) within a unified lattice field theory framework, complementing lattice QCD studies at the quark and gluon level Tuo et al. (2019); Detmold and Murphy ; Davoudi et al. (2024); Feng et al. (2020); Ma et al. (2024b).
Acknowledgements.
We are grateful for helpful discussions on the nuclear interactions with Dean Lee, Ulf-G. Meißner, and Shihang Shen. We appreciate Hermann Krebs for the choice of axial currents, Young-Ho Song for the design of Monte Carlo sampling method, and Jiangming Yao for helpful comments. X.F. and T.W. were supported in part by NSFC of China under Grants No. 12125501, No. 12293060, No. 12293063, and No. 12141501, and by the National Key Research and Development Program of China under Grant No. 2020YFA0406400. B.N.L. was supported by NSAF Grant No.U2330401 and by the National Natural Science Foundation of China under Grant No.12275259.
[…]
Appendix A Nucleon-nucleon Scattering
We use the spherical wall method to calculate nucleon-nucleon phase shifts on the lattice SuppNN1 ; SuppNN2 ; SuppNN3 . The fitting results of $n$-$p$ scattering for the full Hamiltonian $H$ (blue dash-dot line) and the nonperturbative Hamiltonian $H_{0}$ (black dashed line) is shown in Fig. S 1, with the center-of-mass momentum $P\leq 200$ MeV. We also show the phase shifts extracted from the Nijmegen partial wave analysis SuppNN4 , represented by the red points. For $H_{0}$, results are only fitted to $^{1}S_{0}$ and $^{3}S_{1}$ channels, so reasonable agreement with the experimental data appears in these two channels. The results for other channels are predictions based on the $S$-waves fit, and the agreement is not very satisfactory. We could in principle make an improvement by increasing the value $\Lambda^{\prime}_{\pi}$ in Eq. (13) or adding more operators in $H_{0}$, but this would introduce additional sign problems. For $H$, the fitting shows good agreement with the experiment in most channels. In several $D$-wave channels there are some deviations, which can be eliminated if operators beyond N2LO are included.
Figure S 1: $n$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ for $H$ and $H_{0}$. The red points are from Nijmegen partial wave analysis. The blue and black lines denote fitting results of the full Hamiltonian $H$ and the nonperturbative Hamiltonian $H_{0}$, respectively.
Figure S 2: $p$-$p$ scattering phase shifts versus the center-of-mass momentum $P$ in the $^{1}S_{0}$ channel. The black points with errors are from the Nijmegen partial wave analysis. The gray line denotes the fit result.
For the isospin-breaking potentials $V_{pp}$ and $V_{nn}$, we tune the value of $C_{pp}$ to reproduce the $p$-$p$ phase shifts in the $^{1}S_{0}$ channel SuppNN4 . The fitting result is given in Fig. S 2. For $V_{nn}$, we fit it to the experimental value of the $n$-$n$ scattering length SuppNN6 and effective range SuppNN7 in the $^{1}S_{0}$ channel:
$a_{nn}=-18.95\pm 0.40\ \mathrm{fm},\quad r_{nn}=2.75\pm 0.11\ \mathrm{fm}.$ (39)
Appendix B Perturbative Expression of the RGTME
In this section, we give detailed expressions for the perturbative expansion of the RGTME.
We write Eq. (27) more explicitly as
$\mathrm{GT}(t)=\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi_{f}|M^{N_{t}/2}\mathcal{J}M^{N_{t}/2}|\Phi_{i}\rangle}{\sqrt{\langle\Phi_{i}|M^{N_{t}}|\Phi_{i}\rangle}\sqrt{\langle\Phi_{f}|M^{N_{t}}|\Phi_{f}\rangle}}.$ (40)
According to Eq. (28), the full transfer matrix is separated into two parts,
$M=M_{0}+\Delta M.$ (41)
Then Eq. (40) can be expanded as a power series of $\Delta M$:
$\mathrm{GT}(t)=\mathrm{GT}^{(0)}(t)+\mathrm{GT}^{(1)}(t)+\mathcal{O}(\Delta M^{2}).$ (42)
To write down the expressions for $\mathrm{GT}^{(0)}$ and $\mathrm{GT}^{(1)}$, we introduce the notation
$|\Phi_{i/f,0}^{t/2}\rangle=M_{0}^{N_{t}/2}|\Phi_{i/f}\rangle$ (43)
to describe the evolution of trial states under the nonperturbative Hamiltonian $H_{0}$. We also introduce
$|\delta\Phi_{i/f}^{t/2}\rangle=\sum_{k=1}^{N_{t}/2}M_{0}^{k-1}\Delta MM_{0}^{N_{t}/2-k}|\Phi_{i/f}\rangle$ (44)
as the first-order correction to Eq. (43). With Eq. (43) and (44), it easy to derive
$\mathrm{GT}^{(0)}(t)=\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi_{f,0}^{t/2}|\mathcal{J}|\Phi_{i,0}^{t/2}\rangle}{\sqrt{\langle\Phi_{i,0}^{t/2}|\Phi_{i,0}^{t/2}\rangle\langle\Phi_{f,0}^{t/2}|\Phi_{f,0}^{t/2}\rangle}}$ (45)
and
$\mathrm{GT}^{(1)}(t)$ $=$ $\frac{\sqrt{2J_{f}+1}}{g_{A}\langle J_{i}M_{i},11|J_{f}M_{f}\rangle}\frac{\langle\Phi_{f,0}^{t/2}|\mathcal{J}|\delta\Phi_{i}^{t/2}\rangle+\langle\delta\Phi_{f}^{t/2}|\mathcal{J}|\Phi_{i,0}^{t/2}\rangle}{\sqrt{\langle\Phi_{i,0}^{t/2}|\Phi_{i,0}^{t/2}\rangle\langle\Phi_{f,0}^{t/2}|\Phi_{f,0}^{t/2}\rangle}}$ (46)
- $\mathrm{GT}^{(0)}(t)\left[\frac{\langle\Phi_{i,0}^{t/2}|\delta\Phi_{i}^{t/2}\rangle+\langle\delta\Phi_{i}^{t/2}|\Phi_{i,0}^{t/2}\rangle}{2\langle\Phi_{i,0}^{t/2}|\Phi_{i,0}^{t/2}\rangle}+i\leftrightarrow f\right].$
Appendix C More about the Monte Carlo Method
In this work, we need to calculate matrix elements of the following form
$\langle\Psi^{t/2}_{a,0}|\mathcal{O}|\Psi^{t/2}_{b,0}\rangle,\ a,b\in\{i,f\},\ \mathcal{O}\in\{1,\mathcal{J}\},$ (47)
as is shown in Eqs. (45) and (46). Applying auxiliary field transformation, we have,
$\langle\Phi_{a,0}^{t/2}|\mathcal{O}|\Phi_{b,0}^{t/2}\rangle=\int\mathcal{D}s\ e^{-S[s]}Z_{a,b}^{\mathcal{O}}[s].$ (48)
Here $s$ is the auxiliary field located at all the lattice points and $S[s]$ is a Gaussian-type function. The interaction between the nucleons and the auxiliary field is included in $Z_{a,b}^{O}[s]$. After the transformation, both the nominators and the denominators in Eqs. (45) and (46) become integrals over $s$. We evaluate them through generating a set of auxiliary fields which distribute according to a given probability distribution function $P[s]$. A proper choice of $P[s]$ can help improve the sampling efficiency and reduce statistical noises. As both the initial and final states of the nucleus nucleus contribute to the $\beta$-decay process, we take the following form of $P[s]$:
$P[s]\propto c_{i}e^{-S[s]}|Z^{\mathcal{O}=1}_{i,i}[s]|+c_{f}e^{-S[s]}|Z^{\mathcal{O}=1}_{f,f}[s]|,$ (49)
with $c_{i}$ and $c_{f}$ positive numbers. Their values are tuned to suppress statistical errors. The update of the auxiliary field is accelerated by the shuttle algorithm proposed in Ref. Supp2 .
[…]